OpenStudy (anonymous):

3/4% of what number is 8.35? Round to the nearest whole number A.111 B.1113

OpenStudy (mrnood):

what is 3/4 % written as a decimal?

OpenStudy (mrnood):

If you want help - then please respond. You can start by writing 3/4 as a decimal then to make that into a percentage you divide by 100 so 3/4 = ? (as decimal - just type it into a calculator if you don't know) then take that decimal and divide by 100

OpenStudy (anonymous):

3/4 as a decimal is .75

OpenStudy (mrnood):

ok good so divide that by 100 and write as a decimal

OpenStudy (anonymous):

0.0075

OpenStudy (mrnood):

Ok so 3/4% of something is the same a 0.0075 times something if the answer to your question is the number x then x times 0.0075 = 8.35 also written a 0.0075x = 8.35 can you solve what x is from that equation?

OpenStudy (anonymous):

x = 3340/3 which = 1113

OpenStudy (anonymous):

Thank you

OpenStudy (mrnood):

not sure quite how you jumped from my equation to your answer. But well done anyhow.